OpenStudy (anonymous):

what is the area of 5 5 9 9

OpenStudy (anonymous):

is there a figure for that?

OpenStudy (anonymous):

5* 9 would be the answer

OpenStudy (anonymous):

its a rectangle @jackmullen55

OpenStudy (anonymous):

It could also be a trapezoid

OpenStudy (anonymous):

it is a rectangle

OpenStudy (anonymous):

ok so 9*5 is........?

OpenStudy (anonymous):

DONT GIVE THE ANSWER JACK!

OpenStudy (anonymous):

Oops damn tourettes

OpenStudy (anonymous):

huh?

OpenStudy (anonymous):

never mind just looked up what it is @jackmullen55 its a really intresting thing this tourettes

OpenStudy (kitkat16):

A=w(l)

OpenStudy (anonymous):

Yes it really is. It sucks to have to 9*5 = 45. Damn. There it goes again... No seriously I do actually have tourettes

OpenStudy (anonymous):

yeah what @KitKat16 said. its fine jack

OpenStudy (anonymous):

sory my thing would not tipe

OpenStudy (kitkat16):

a=5(9) and it is obvious that it is a rectangle. lol

OpenStudy (anonymous):

the answer is b thanks boys ;)

OpenStudy (anonymous):

i aint a guy tho! thats insulting!

OpenStudy (anonymous):

Well there is no indication of gender. boys, guys, etc. is typically just a general thing

OpenStudy (denisaboichuk):

a=5(9)